Session: Nov 2004 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Xinyong Huang at Thu Oct 28 22:49:30 2004) Success :
The label numbers of the discs are: SHAO-015 and SHAO-007. Please produce the codes, bcuase of we haven't the code scanner.

Medicina (Zacchiroli at Fri Oct 29 00:16:42 2004) Success :

Urumqi (Liu Xiang at Fri Oct 29 03:03:06 2004) Success :
It is snow and rain in beginning of experiment, then a few hours after it is cloudy and clear sooner. UT20:36--58 control PC dead. Two disk-packs were used.

Torun (K. Borkowski at Fri Oct 29 05:49:05 2004) Success :
No known problems. Three packs recorded.

Onsala (Michael Lindqvist at Fri Oct 29 06:44:36 2004) Minor failures :
After about 15min of recording we had problem with the disk module (JOD-0001). We had to reboot the Mark 5 unit. Problems with the noisecal (i.e, bad tsys-values) during some scans. It did not fire. Can be seen if one compare the tpi/tpical-values where the tsys-valus are bad.

Effelsberg (D.Graham at Fri Oct 29 08:27:50 2004) Minor failures :
Tsys measurements poor before 14:50 as they are all made while slewing, otherwise OK

Noto (P. Cassaro at Fri Oct 29 09:18:18 2004) Success :

Urumqi (Liu Xiang at Fri Oct 29 13:18:45 2004) Success :
Sorry, we forgot to put 'red label' on disk-packs, which are UAO-0012, UAO-0014

Jodrell_Bank (Alastair Gunn at Mon Nov 1 01:06:23 2004) Minor failures : SLEW
The Lovell Telescope was used. Sorry, had to edit this schedule by hand to comply with Lovell slewing restrictions. Remained on primary target for many of the calibrator scans. Connection to Mk5 lost between 23:31 and 23:41 UT. No other known problems.

Westerbork (Tony Foley at Mon Nov 1 11:12:04 2004) Success :
Wb used in phased array mode. Mark5A gave an error code 6 about 16:00UT - but seems OK otherwise